OpenStudy (anonymous):

A ladder 13 meters long rests on horizontal ground and leans against a vertical wall. The foot of the ladder is pulled away from the wall at the rate of 0.6 m/sec. How fast is the top sliding down the wall when the foot of the ladder is 5 m from the wall?

OpenStudy (anonymous):

This is the one where I can't figure out the equation.

OpenStudy (anonymous):

Let the horizontal distance from the wall be x and let the height of the top of the ladder be y. Note that the length of the ladder is the hypotenuse of a right triangle, so we have: \[ x^2 + y^2 = 13^2 = 169. \]The statement about the rate talks about a derivative (in specific, dx/dt - because it is the rate at which x is being change, wrt t). Do you see where it's heading or would you like some more guidance?

OpenStudy (anonymous):

so, I just take the derivative and solve for y'?

OpenStudy (anonymous):

2x*x'+2y*y'=0?

OpenStudy (anonymous):

plug in x, y and x'

OpenStudy (anonymous):

Wait, I'm confused again.

OpenStudy (anonymous):

I don't have a y

OpenStudy (anonymous):

or am I taking the derivative in respect to y?

OpenStudy (anonymous):

even then, I'm confused.

OpenStudy (anonymous):

Interpret the question. What is it asking you - as a derivative? I'm trying not to give away the solution so that you can think about the problem a little. Bear with me.

OpenStudy (anonymous):

Ok, no worries. Um..

OpenStudy (anonymous):

I'm assuming its in respect to time.

OpenStudy (anonymous):

but when I get the derivative, I don't know what to plug in for y.

OpenStudy (anonymous):

I know to plug in 0.6 to x' and 5 to x and I know I'm looking for y'

OpenStudy (anonymous):

Is that wrong?

OpenStudy (anonymous):

I figured it out I think.

OpenStudy (anonymous):

I just get y by plugging in x to the original equation. I feel silly.

OpenStudy (anonymous):

5, 12, 13

OpenStudy (anonymous):

right.

OpenStudy (radar):

You can solve for y when x = 5 use pythgorean like Tonks did.

OpenStudy (radar):

Don't be surprised if dy/dt is negative.

OpenStudy (anonymous):

Ok. Thanks!